### v2

# Two-phase formation of galaxies: the coevolution between galaxies and dark matter halos

Abstract
We use FIRE-2 cosmological zoom-in hydrodynamic simulations to investigate the co-evolution between Milky Way-size galaxies and their host dark matter halos. We find that the formation of these galaxies follows a two-phase pattern, with an early phase featured by hot dynamics, bulge-dominated structure and bursty star formation, and a later phase featured by cold dynamics, disk-dominated structure and steady star formation. The transition times of these galaxy properties are tightly correlated with the time when the host halo transits from fast to slow accretion, suggesting that the two-phase assembly of halos drive the two-phase formation of galaxies. The physical origin of dynamical hotness can be summarized into two modes of star formation: a scattered mode in which stars form at large radii within cold gas streams associated with fast assembly of halos, and a concentrated mode in which stars form at small radii through violent fragmentation from globally self-gravitated gas when halo assembly is about to slow down. Cold gaseous and stellar disks can form only when the conditions of the two modes are removed by the stall of halo assembly and the reduction of gas by feedback processes. The two modes of star formation leave distinct imprints on the structural properties of high-redshift galaxies, providing implications to be tested by JWST and future observations.
keywords:
galaxies: kinematics and dynamics – galaxies: formation – galaxies: haloes – galaxies: high-redshift
pubyear: 2025pagerange: Two-phase formation of galaxies: the coevolution between galaxies and dark matter halosA
1 Introduction
Galaxies exhibit diverse morphologies ranging from those dominated by bulges to those dominated by disks – known as the Hubble sequence (e.g. Hubble, 1926, 1936; Sandage, 1961; Kormendy, 1979; Conselice, 2014). Observations with resolved kinematic measurements suggest that the morphologies of galaxies are closely related to the motion of stars, with bulges being populated mainly by stars with random orbits and disks with ordered rotation (e.g. Cappellari, 2016; Zhu et al., 2018). Historically, these two types of motion were inferred to reflect the evolution stages of galaxies, with ordered rotation being the early products of galaxy formation owing to the collapse of gas with initial angular momentum in halos, and random motion being generated by later ‘heating’ processes such as mergers (e.g. Hernquist, 1992; Shen et al., 2003; Lacey et al., 2016). Such a picture of galaxy formation and morphological transformation has been widely adopted in theoretical models and achieved great success in reproducing the observed properties of the galaxy population observed in the local Universe, such as the size-mass relations (Shen et al., 2003; van Dokkum et al., 2008, 2015; Bezanson et al., 2009; Newman et al., 2012; Graham, 2013; van der Wel et al., 2014a; Somerville et al., 2018; Miller et al., 2019) and the bimodality in the color-stellar mass relation (Strateva et al., 2001; Baldry et al., 2004).
[…]
Along this line, Mo et al. (2024, hereafter MCW24) proposed a ‘two-phase’ scenario of galaxy formation, in which the two morphological components of galaxies, bulge and disk, naturally arise from cosmological structure formation. This scenario is motivated by the fact that the assembly of individual CDM halos follows a general pattern consisting of two phases: an early phase of fast accretion and a later phase of slow accretion – a discovery made decades ago in N-body simulations (Zhao et al. 2003; see also Wechsler et al. 2002; Zhao et al. 2009; Correa et al. 2015). During the fast phase of a halo, frequent perturbation of gas is produced by the rapid change of gravitational potential associated with the fast accretion, and is subsequently amplified by shocks in supersonic turbulence generated by rapid gas inflow. Dense sub-clouds are produced by the fragmentation of the collapsing gas. The rich amount of gas and the strong turbulence brought in by the fast accretion imply that the fragmentation starts once the gas becomes self-gravitating and that the formation of dense sub-clouds proceeds rapidly before the gas can dissipate to form a dynamically cold disk. A galaxy formed in this phase is thus manifested as a self-gravitating gas cloud (SGC) filled with sub-clouds with random orbits, with star clusters born out of the sub-clouds inheriting their hot dynamics. Dynamically cold gaseous and stellar disk can form only after the host halo transits to the slow phase and the gas content is reduced by feedback processes. Implemented semi-analytically into subhalo mergers trees, this scenario can reproduce a wide range of observables, and coherently predict the co-evolution among halos, galaxies, supermassive black holes (SMBHs), and globular clusters (GCs) from the seeding era to the present day (see Mo et al., 2024; Chen et al., 2024, 2025b, 2025a, for details).
One important prediction of the two-phase scenario is a counterintuitive order in the build-up of the Hubble sequence: the dynamically hot (bulge) component of a galaxy is actually built first, and followed by the dynamically cold (disk) component, with a transition corresponding to that in halo assembly from fast to slow. This scenario seems to be in conflict with the historical picture of galaxy formation and morphological transformation mentioned above, but becomes increasingly supported by recent hydrodynamic simulations (Dekel et al., 2009; Pillepich et al., 2019; Hopkins et al., 2018; Jung et al., 2025) and observations (Kassin et al., 2012; Swinbank et al., 2012; Wisnioski et al., 2015; Simons et al., 2017; Price et al., 2020). However, many questions remain unanswered. For example, it is unclear whether the fast accretion can indeed sustain a gas-rich and turbulent SGC, whether sub-clouds can actually form in the SGC quickly before settling into a rotation-supported disk, and whether feedback processes are strong enough to reduce the gas content to allow the formation of a stable disk. Given the challenges in obtaining a complete sample of high-redshift galaxies to perform dynamical decomposition, it is still difficult to answer these questions directly using observational data. On the other hand, modern zoom-in hydro simulations with cosmological initial conditions are able to resolve key processes involved in the formation of individual galaxies, thus providing opportunities to test the viability of the two-phase model.
In this paper, we use a set of such simulations performed as a part of the Feedback In Realistic Environments (FIRE) project (Hopkins et al., 2018) to analyze the co-evolution between galaxies and dark matter halos, focusing on the aforementioned questions. This paper is organized as follows. In §2, we describe the simulation and the method to quantify halos and galaxies. In §3, we verify the two-phase growth pattern of galaxies and its relation to the two-phase assembly of halos. In §4, we analyze in detail the physical processes that drive the two-phase growth of galaxies. In §5, we provide implications for observations in a search for evidence induced by these processes. Finally, we summarize and discuss our findings in §6.
2 The Simulation
Table 1: The sample of FIRE-2 galaxies used in this paper. The properties listed in the columns are: (i) the name of the galaxy, with the reference first introducing the galaxy indicated in the parentheses (A, Garrison-Kimmel et al. 2019a; B, Garrison-Kimmel et al. 2019b; C, Garrison-Kimmel et al. 2017; D, Wetzel et al. 2016; E, Hopkins et al. 2018; F, Samuel et al. 2020); (ii) $M_{\mathrm{vir},\mathrm{dm}}$, the mass of dark matter enclosed in the virial radius of the host halo at $z=0$; (iii) $M_{*}$, stellar mass of the galaxy; (iv) $R_{50,*}$, half-stellar-mass radius using all star particles linked to the galaxy; (v) $R_{\rm 50,bulge}$, half-stellar-mass radius using only bulge stars; (vi) $R_{90,*}$, the radius enclosing 90% of the stellar mass, using all star particles linked to the galaxy; (vii) $t_{\mathcal{B}}$, the transition time of burstiness; (viii) $t^{(\mathrm{start})}_{\mathcal{H}}$$t^{(\mathrm{end})}_{\mathcal{H}}$, the transitional interval of hotness; (ix) $^{(\mathrm{start})}_{\mathrm{thin}}tt$$^{(\mathrm{end})}_{\mathrm{thin}}$, the transitional interval of thin-disk fraction; (x) $t_{\rm bulge,1/2}$, the half-mass formation time of stellar bulge; (xi) $t^{(\mathrm{start})}_{\gamma}$$t^{(\mathrm{end})}_{\gamma}$, the transitional interval of halo accretion rate; (xii) $t_{\rm h,1/2}$, the half-mass formation time of the host halo; (xiii) $t_{\text{h, Vpeak}}$, the epoch when the maximum circular velocity of the halo reaches the peak during the assembly history. See §2 for the details of the sample and the definitions of the properties.
Name (ref) $_{\mathrm{vir},\mathrm{dm}} MM_{*}$ $R_{\rm 50,*}$ $R_{\rm 50,bulge}$ $R_{\rm 90,*}$ $t_{\mathcal{B}}$ $t^{(\mathrm{start})}_{\mathcal{H}}$$^{(\mathrm{end})}_{\mathcal{H}} tt$$^{(\mathrm{start})}_{\mathrm{thin}}$$t^{(\mathrm{end})}_{\mathrm{thin}}$ $t_{\rm bulge,1/2}$ $t^{(\mathrm{start})}_{\gamma}$$t^{(\mathrm{end})}_{\gamma}$ $t_{\rm h,1/2}$ $t_{\text{h, Vpeak}}$
$\rm M$$_{\odot}$] $\rm M$$_{\odot}$] $\rm kpc$] $\rm kpc$] $\rm kpc$] $\rm Gyr$}] $\rm Gyr$}] $\rm Gyr$}] $\rm Gyr$}] $\rm Gyr$}] $\rm Gyr$}] $\rm Gyr$}]
[…]
In this paper, we use the cosmological hydrodynamic zoom-in simulations from FIRE-2, a part of the FIRE project, performed using a multimethod gravity plus (magneto)hydrodynamics code GIZMO (Hopkins, 2015), with updated numerical methods and physical models as detailed in Hopkins et al. (2018). FIRE-2 adopted the mesh-free Lagrangian Godunov (MFM) method for hydrodynamics, which provides adaptive spatial resolution while maintaining exact conservation of mass, energy, and momentum, excellent angular momentum conservation, and accurate shock capturing. Radiative heating and cooling for gas across a temperature range of $\sim 10$$10^{10}\,{\rm K}$ are included, involving free-free, photoionization/recombination, Compton, photo-electric, metal-line, molecular, fine-structure, dust collisional, and cosmic-ray processes. Eleven elements (H, He, C, N, O, Ne, Mg, Si, S, Ca, Fe) are self-consistently tracked. Unresolved turbulent diffusion of metals in the gas is explicitly modeled (Hopkins, 2016; Su et al., 2017; Escala et al., 2018). The simulations also include photoionization and heating from a redshift-dependent, spatially uniform ultraviolet (UV) background (Faucher-Giguère et al., 2009) that reionizes the universe at $z\sim 10$. Collisionless star particles, representing ensembles of stars, are bred from molecular gas satisfying self-gravitating, self-shielding, Jeans unstable conditions and exceeding a minimum density threshold of $1000\,{\rm cm}^{-3}$ (Krumholz & Gnedin, 2011). FIRE-2 implemented distinct channels of stellar feedback, including that from OB stars, AGB mass-loss, type Ia and type II supernovae, photoelectric heating, and radiation pressure (Hopkins et al., 2018, 2020). Such feedback regulates the star formation efficiency of star-forming gas to a level of $\sim 1\%$$10\%$ per free-fall time (Faucher-Giguère et al., 2013; Hopkins, 2017; Orr et al., 2018). We note that feedback from accreting SMBHs is not included in the simulations.
[…]
Figure 1: Visualization and profiles of the galaxy m12f at $z=0$. (a)(d), 2-D stellar mass surface density, viewed edge-on (row 1) and face-on (row 2). The four columns show the results for all stellar particles, and those belonging thin disk ($\epsilon\geq 0.8$), thick disk ($0.2\leq\epsilon<0.8$) and bulge ($\epsilon<0.2$), respectively. (e), mass-weighted distribution of circularity ($\epsilon$) of star particles. Colored solid curve shows the distribution of all star particles within $R_{\rm 90,*}$, with three colors indicating the three kinematic components, respectively. Black solid curve shows the distribution of only young star particles (age $100\,{\rm Myr}$) within $R_{\rm 90,*}$. The mass fraction of each component in indicated in the legend for all/young stars. (f), mass surface density profiles, obtained using all (solid curves) and young (dashed curves) stars, showing in total (black) and separately for the three kinematic components (colored). (g), hotness as a function of radius, obtained by using all (solid curve) and young (dashed curve) stars enclosed within the radius. We mark the galaxy sizes, $R_{\rm 50,*}$ (blue) and $R_{\rm 90,*}$ (red), by circles in (a2) and vertical lines in (f) and (g).
Figure 2: Properties of the galaxy m12f and its host halo as functions of the cosmic time ($t$) or redshift ($z$). Each panel shows a different property. (a), stellar mass in total (gray) and separately for the three kinematic components (colored), and $M_{\rm vir}$ of the host halo (black). Three vertical lines indicate the epochs at which the bulge mass reaches 20%, 50% and 80%, respectively, of its final value at $z=0$, with that corresponding to 50% denoted as $t_{\rm bulge,1/2}$. (b), the growth rate of the three kinematic components (colored). For reference, we show the accretion rate of baryons, defined as the growth rate of $M_{\rm vir}$ times the cosmic baryon fraction (black). (c), mass fractions of the three kinematic components, using young stars only. Grey vertical band indicates the transitional interval, $_{\rm tnin}^{\rm(start)}tt$$_{\rm thin}^{\rm(end)}$, of the thin-disk fraction. (d), dynamical hotness of the galaxy ($\mathcal{H}$), calculated using young stars only. Grey vertical band indicates the transitional interval, $t_{\mathcal{H}}^{\rm(start)}$$t_{\mathcal{H}}^{\rm(end)}$ of $\mathcal{H}$. (e), burstiness of star formation ($\mathcal{B}$). Vertical line indicates $t_{\mathcal{B}}$, the epoch when the star formation transits from bursty to steady (defined by $\mathcal{B}=0.2$, indicated by the dashed horizontal line). (f), cold-gas spin $\lambda_{\text{cold-gas}}$, and fractions of cold ($f_{\text{cold-gas}}$), warm and hot gas. (g), (h), virial velocity ($V_{\rm vir}$) and specific growth rate of the host halo. Grey vertical bands indicate the transitional interval, $t_{\gamma}^{\rm(start)}$$t_{\gamma}^{\rm(end)}$, of halo assembly, defined by $\gamma=3/8$0 (horizontal lines in h). In (c), (d), (g) and (h), each solid curve is obtained from the simulation, while the associated dashed curve is obtained by a parametric fitting (see §2.1 and §2.2 for details). This figure clearly demonstrate the two-phase evolution of the galaxy and its host halo. See §3 for details.
2.1 Properties of the simulated galaxies
Table 2: Properties of galaxies and halos defined and analyzed in this paper.
Property Description
$\epsilon$ Orbital circular of a particle (§2.1.1)
$f_{\rm thin}$ Mass fraction of thin disk (§2.1.1); similar definitions are made for thick disk ($f_{\rm thick}$) and bulge ($f_{\rm bulge}$)
$\mathcal{H}$ Dynamical hotness (§2.1.2)
$\mathcal{B}$ Burstiness of the star formation (§2.1.3)
$n$ Sérsic index of bulge (§2.1.4)
$c/a$ Minor to major axis ratio of bulge (§2.1.4); similar definition is made for intermediate to major axis ratio ($b/a$)
$T$ Triaxiality of bulge (§2.1.4)
$f_{\lambda}$ Ratio between cold-gas fraction and cold-gas spin (§2.1.5)
$f_{\text{gas+star}}$ Mass fraction of cold gas plus young stars within some radius (§2.1.5)
$\gamma$ Specific growth rate of halo (§2.2)
2.1.1 Stellar kinematics and decomposition
We follow the decomposition method of Abadi et al. (2003) to separate member star particles of each galaxy at a given snapshot into three components: bulge, thick disk and thin disk. The method is based on the circularity of each star particle,
$\epsilon=j_{z}/j_{\rm c}(E)\,$, (1)
where $j_{z}$ is the angular momentum of that particle projected to the direction of $\bm{J}_{\rm tot}$, the net angular momentum of all stellar particles within $R_{\rm 90,*}$ at that snapshot (Gurvich et al., 2023), $j_{\rm c}(E)$ is the angular momentum of a circular orbit with the same energy ($E$) as that of the star particle. Particles with $1\leqslant\epsilon<0.2$, $0.2\leqslant\epsilon<0.8$ and $0.8\leqslant\epsilon\leqslant 1$ are categorized as bulge, thick disk and thin disk, respectively. Previous studies adopted more rigorous kinematic classification methods (e.g. Du et al., 2019, 2020; Zana et al., 2022; Liang et al., 2025), such as classifications in the plane of angular momentum versus orbital energy. However, our work focuses primarily on morphological transitions, particularly in the evolution of the significance of the disk component, which can be characterized robustly using the angular momentum alone. We therefore employ a simplified kinematic decomposition based solely on the angular momentum to define the three components. We define the direction of $\bm{J}_{\rm tot}$ as the normal direction (the $z$ axis) of the disk ($x$-$y$) plane. The projections of a simulated galaxy perpendicular to and along the $z$-axis are defined as the edge-on and face-on views of the galaxy, respectively.
Fig. 1 shows an example galaxy, m12f, at $z=0$, with the three components shown separately and in total, viewed edge-on and face-on. Here, we include all particles within $R_{\rm 90,*}$. Panel (e) shows the distribution of $\epsilon$ for all stellar particles within $R_{\rm 90,*}$, weighted by their masses. Panel (f) shows the face-on surface density profile of the three components and of the total, using all particles within $R_{\rm 90,*}$. The thin disk is extended, dominates the surface density in the outer region, and follows an exponential profile at $r\gtrsim 3\,{\rm{kpc}}\approx R_{\rm 50,*}$. In the inner region ($r\lesssim 1\,{\rm{kpc}}$), however, the thin disk is truncated. This truncation is similar to that found in other simulations (Du et al. 2020, see their fig. 1; Buck et al. 2019, see their fig. 2) and in observations with spatially-resolved spectroscopy (Zhu et al. 2018, see their fig. 8 and §4), and suggests a potential caveat of using an untruncated profile to fit the entire cold-disk component. On the other hand, the thick disc is more vertically extended, reflecting its hotter dynamics with a higher degree of random-motion support for the vertical structure. Both the images (c1, c2) and the surface density profile (yellow in f) suggest the existence of an inner, concentrated component at around $1\,{\rm{kpc}}$, in alignment with the concept of slow-rotating ‘pseudo-bulge’ in the literature (Kormendy & Kennicutt, 2004). The bulge dominates the inner $0.5\,{\rm{kpc}}$, but it also has a faint extension to the outer region ($\gtrsim 10\,{\rm{kpc}}$), consistent with being a combination of the ‘classical bulge’ and the stellar halo found in the literature (Genel et al., 2015; Du et al., 2020; Zana et al., 2022). The three components have similar mass within $R_{\rm 90,*}$, as can be seen from the mass-weighted distribution of $\epsilon$ (panel e), consistent with the results of an earlier analysis of the FIRE-2 by Yu et al. (2021). All these show that FIRE-2 is able to produce realistic kinematics of MW-size galaxies, with components similar to those found in observations.
To quantify the kinematics of star-forming gas within a target galaxy, we utilize in-situ young stars, defined as bound stellar particles with ages $\leq 100\,{\rm Myr}$ and birthplaces within $\leq R_{90,*}$ from the center of the galaxy. This selection is made to avoid complexity introduced by potential post-formation heating of old stars and the contamination of ex-situ stars. Hereafter, we use these in-situ young stars to define transition times in the morphological type of galaxy. As shown in Fig. 1(e), the distribution of $\epsilon$ of in-situ young stars at $z=0$ is sharply peaked around $\epsilon=1$. Most ($\approx 93\%$) of the in-situ young stars at this epoch belong to the thin disk, and only a small fraction ($\approx 1.6\%$) is categorized into the bulge component. Fig. 1(f) reinforces this conclusion by comparing the face-on surface densities of in-situ young stars (dashed lines) with those of all stars (solid lines), which shows that in-situ young stars are mostly contained in the extended thin disk.
Fig. 2(a) shows the stellar mass as a function of the cosmic time ($t$), for m12f as a total and separately for the three components. Here, we first measure the radius that encloses 90% of stellar mass of the corresponding component and then estimate the stellar mass within this radius. For reference, we also include the mass of the host halo, $M_{\rm vir}$. In parallel, panels (b) and (c) show the corresponding mass growth rate, and mass fraction of in-situ young stars in different kinematic components, respectively. Remarkably, at $z\gtrsim 1$, stars are mostly born in dynamically hot (bulge and thick-disk) components, and the fraction in the cold, thin disk component is very small. The mass fractions in the bulge and thick-disk components are comparable and show similar redshift dependence. At $z\approx 1$, the star-formation activity rapidly migrates to the thin disk, outpacing that in the hot components within a timescale of $\approx 1\,{\rm Gyr}$. Thereafter, the mass in the bulge changes little, while that in the thin disk increases steadily. The behavior of $M_{\rm bulge}$ versus $t$ prompts the definition of a formation time, $t_{\rm bulge,1/2}$, which we set to be the time when the bulge has reached half ($50\%$) of its final mass and is marked by a vertical line in panel (a). For reference, we also mark the times for the formation of $20\%$ and $80\%$ of the final bulge mass.
The clear transition of star formation from the dynamically hot to cold components shown in Fig. 2(c) motivates us to define a transition time to separate the early, hot phase from the late, cold phase, for each individual galaxy. To do this, we fit the fraction of young stars in the thin-disk component ($f_{\rm thin}$) as a function of the cosmic time ($t$), as shown by the blue solid curve in Fig. 2(c) for example, by the following piecewise function:
[…]
Here, $f_{1}$ and $f_{2}$ are the initial and final values of $f_{\rm thin}$, respectively, introduced to describe the shape of $f_{\rm thin}$ seen in Fig. 2(c) at $t<t_{1}$ and $t\geqslant t_{2}$. The change of $f_{\rm thin}$ in the transitional interval between $t_{1}$ and $t_{2}$ is described by a linear function. The fitting result for m12f is shown by the blue dashed curve in Fig. 2(c), with the duration of the transitional phase ($z\approx 0.5$11, or $t\approx 5.5$$8.5\,{\rm Gyr}$) indicated by the grey shade. To avoid ambiguity, we will denote $t_{1}$ and $t_{2}$ as $t_{\rm thin}^{\rm(start)}$ and $t_{\rm thin}^{\rm(end)}$, respectively. We use the mean value of the two,
$t_{\rm thin}=\frac{1}{2}\left[t_{\rm thin}^{\rm(start)}+t_{\rm thin}^{\rm(end)}\right]\,$, (3)
to represent the epoch of the morphology transition, and use the interval $_{\rm thin}^{\rm(start)}, tt$$_{\rm thin}^{\rm(end)}$] to bracket the uncertainty. For comparison, we denote the fraction in the thick-disk and bulge components as $f_{\rm thick}$ and $f_{\rm bulge}$, respectively.
2.1.2 Dynamical hotness and phases of galaxies
To describe the transition behavior of FIRE-2 galaxies while avoiding arbitrariness in the threshold separating the bulge and disk components, we define another physical parameter, hotness, as the ratio between the stellar velocity dispersion and the total velocity within some radius, $R$:
$\mathcal{H}=\frac{\sigma_{\rm v}}{\left|\bm{v}\right|}\,.$ (4)
Here, the total velocity is computed as the mass-weighted total velocity of star particles:
$\left|\bm{v}\right|=\sqrt{\frac{\sum_{i}M_{i}{\bm{v}}_{i}^{2}}{\sum_{i}M_{i}}}\,$, (5)
where $M_{i}$ and ${\bm{v}}_{i}$ are the mass and velocity vector, respectively, of the $i$-th star particle, and the summation is over all in-situ young star particles within $R$ that are bound to the galaxy.
To remove rotational motion from the computation of $\sigma_{\rm v}$, we find the azimuthal velocity, $v_{\phi}$, of each star particle in the cylindrical coordinate system (centered on the galaxy center, with the cylindrical axis parallel to $\bm{J}_{\rm tot}$). We obtain the radial profile, $v_{\phi}(r_{\rm p})$, as the mass-weighted $v_{\phi}$ of the star particles in bins of cylindrical radius, $r_{\rm p}$. We define the rotation-subtracted velocity of the $i$-th star particle as ${\bm{v}}^{\prime}_{i}\equiv{\bm{v}}_{i}-v_{\phi}(r_{\rm p})\hat{{\bm{e}}}_{\phi,i}$, with $\hat{{\bm{e}}}_{\phi,i}$ being the unit vector in the azimuthal direction at the location of the star particle. The rotation-subtracted 3-D velocity dispersion is then computed as
$\sigma_{\rm v}=\sqrt{\frac{\sum_{i}M_{i}\left|{\bm{v}}^{\prime}_{i}-\left<{\bm{v}}^{\prime}\right>\right|^{2}}{\sum_{i}M_{i}}}\,$, (6)
where $\left<{\bm{v}}^{\prime}\right$> is the mass-weighted mean value of the rotation-subtracted velocities of star particles. The calculations of $_{\phi}(r_{\rm p}), 𝒗\left<{\bm{v}}^{\prime}\right$> and $\sigma_{\rm v}$ are performed by using the same set of star particles as used for $\left|\bm{v}\right|$. So defined, the hotness $\mathcal{H}$ is a dimensionless quantity, whose square is the fraction of kinetic energy in the form of random motion. A value of $\mathcal{H}\sim 1$ thus corresponds to a dynamically hot system supported purely by random motion, usually a bulge, while $\mathcal{H}=0$ corresponds to a dynamically cold system supported purely by rotation, usually a thin disk.
Fig. 1(g) shows the hotness as a function of radius for m12f at $z=0$ for all stars (solid) and in-situ young stars (dashed). For all stars, $\mathcal{H}$ decreases monotonically with radius. The inner region ($R\lesssim 0.3\,{\rm{kpc}}$) is dispersion-dominated, with hotness $\mathcal{H}\approx 1$. In contrast, on large radii ($R\approx 10\,{\rm{kpc}}$), $\mathcal{H}\approx 0.8$, indicating that random motion is still significant in the outer region, but less dominating compared to that in the inner region. The in-situ young stars have significantly lower hotness, consistent with their lower fraction in hot components (bulge and thick disk), as shown in Fig. 1(e), (f). Prominently, $\mathcal{H}$ of in-situ young stars on large radii of $R\gtrsim 3\,{\rm{kpc}}$ is as low as about $0.3$, indicating that most of the kinetic energy is in the form of ordered rotation, consistent with the dominance of the thin-disk component there (panel f). All these findings support the hotness $\mathcal{H}$ as a viable parameter to quantify the dynamical state of a galaxy within a given radius. In the following, we use $\mathcal{H}$ within $R=R_{\rm 90,*}$ to represent the hotness of a galaxy as a whole.
Fig. 2(d) shows the evolution of the hotness for m12f. The value of $\mathcal{H}$ remains at a constant value of $0.9$ at $z\gtrsim 1$, rapidly decreases during the period of $z\approx 0.5$11, and remains $\approx 0.3$ thereafter. The steps at both high and low $z$, as well as the rapid decrease between them, are similar to those seen in $f_{\rm thin}$ described in §2.1.1 and shown in Fig. 2(c). We therefore use the same piecewise function form as given by Eq. (2) to fit the hotness $\mathcal{H}$ as a function of time. We define the starting and ending times of the transitional stage as $t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(start)}$ and $t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(end)}$, respectively, and use $t_{\rm\scriptscriptstyle\mathcal{H}}$ to denote the mean of them. Dashed curve in Fig. 2(d) shows the fitting result, together with the two transition times. Because of the similarity between $\mathcal{H}(t)$ and $f_{\rm thin}(t)$, the two transition times are also about the same between the two functions.
Based on the shape of $\mathcal{H}(t)$ and the corresponding change in the galaxy kinematics, we define, for each galaxy, the following dynamical phases:
  1. 1.
    The hot phase, defined by the time interval $t<t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(start)}$, in which $\mathcal{H}$ is constantly high;
  2. 2.
    The cold phase, defined by the time interval $t\geqslant t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(end)}$, in which the hotness is constantly low;
  3. 3.
    The transitional phase, defined by the time interval $t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(start)}\leqslant t<t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(end)}$, in which the hotness decreases rapidly with time.
Note that such an evolution pattern was also found by McCluskey et al. 2024, who focused on disk settling. The terms ‘Pre-Disc’, ‘Early-Disc’ and ‘Late-Disc’ they introduced are similar to our definitions of different phases based on $f_{\rm thin}(t)$ instead of $\mathcal{H}(t)$.
2.1.3 Burstiness in star formation
MW-size galaxies in FIRE-2 are found to show bursty (steady) star-formation history in the early (late) time, with transition of the ‘burstiness’ coinciding with the dynamical transition (Stern et al., 2020; Flores Velázquez et al., 2021; Yu et al., 2021, 2023; Hopkins et al., 2023). Following Yu et al. 2021, we define the burstiness of star formation, $\mathcal{B}$, as the ratio of the standard deviation of the instantaneous star formation rate (SFR) within a time window $\Delta t=500\,{\rm Myr}$ to the averaged SFR within the same window:
$\mathcal{B}=\frac{\sqrt{{\rm Var}_{\Delta t}\left[{\rm SFR}\right]}}{\left<{\rm SFR}\right>_{\Delta t}}\,.$ (7)
[…]
2.1.4 Structure of stellar bulge
A particular focus of this paper is the origin and evolution of dynamically hot stellar components (bulges) of galaxies. In the following, we define a number of parameters to quantify the structural properties of stellar bulges. The results will be presented in §5.
To quantify the anisotropy of stellar distribution in a bulge, we approximate the bulge by an ellipsoid whose orientation and axis lengths are defined, respectively, by the eigenvectors and eigenvalues of the weighted inertia tensor,
$\mathcal{I}_{ij}\equiv\frac{\sum_{n}m_{n}x_{n,i}x_{n,j}}{\sum_{n}m_{n}r_{n}^{2}}\,.$ (8)
Here i,j=1, 2, 3i,\,j=1,\,2,\,3 are the three spatial dimensions; $m_{n}$ is the mass of the $n$-th member star particle, and $x_{n,i}$ and $x_{n,j}$ are its galactocentric coordinates in the corresponding dimensions. The elliptical distance of the stellar particle, $r_{n}$, is defined as
$r_{n}=\sqrt{x_{n,1}^{2}+\frac{x_{n,2}^{2}}{q^{2}}+\frac{x_{n,3}^{2}}{s^{2}}}\,$, (9)
where $q\equiv b/a$ and $s\equiv c/a$ are axis ratios; $a$, $b$, $c$ $\propto$ $\sqrt{\lambda_{1}}$, $\sqrt{\lambda_{2}}$, $\sqrt{\lambda_{3}}$ are the lengths of the three axes, respectively; $\lambda_{1}\geqslant\lambda_{2}\geqslant\lambda_{3}$ are the eigenvalues of the inertia tensor $\mathcal{I}_{ij}$. The weighting $r_{n}^{-2}$ is adopted to avoid over-weighting particles at large radii, as introduced by Allgood et al. (2006). The summation is over all star particles within $R_{90,*}$, and the calculations are performed only when the number of particles is greater than 1,000. The eigenvalue-dependence in the definition of the inertia tensor leads to a dependency loop. We thus start from a sphere (i.e. $q=1$ and $s=1$), iteratively compute the inertia tensor and obtain its eigenvectors, and rotate the coordinates $(x_{n,1},x_{n,2},x_{n,3})$ to the frame defined by the eigenvectors, until convergence is reached. Finally, we define the triaxiality parameter,
$T\equiv\frac{1-q^{2}}{1-s^{2}}=\frac{a^{2}-b^{2}}{a^{2}-c^{2}}\,$, (10)
to indicate the anisotropy in the spatial distribution of stars. So defined, a perfect prolate ellipsoid has $T=1$, a perfect oblate ellipsoid has $T=0$, and other cases all have $0<T<1$.
To quantify the concentration of the stellar distribution in the bulge, we project all bulge stars within $R_{\rm 90,*}$ along ${\bm{J}}_{\rm tot}$, measure the surface density profile as a function of the projected galactocentric distance, $r_{\rm p}$, and fit the profile using a Sérsic function:
$\Sigma_{*}(r_{\rm p})=\Sigma_{\rm e,*}\exp\left\{-b_{\rm s}\left[\left(\frac{r_{\rm p}}{R_{\rm e,*}}\right)^{1/n}-1\right]\right\}\,$, (11)
where $R_{\rm e,*}$ is the effective radius, $\Sigma_{\rm e,*}$ is the surface density at that radius, $n$ is the Sérsic index that sets the concentration of the profile, and $b_{\rm s}$ is a parameter introduced to support the meaning of other parameters. The fitting is done by minimizing the $\chi^{2}$ loss function between the function and the simulated profile:
$\chi^{2}(\Sigma_{\rm e,*},b_{\rm s},R_{\rm e,*},n)=\sum_{i}\frac{\left[M_{i}-\hat{M}_{i}(\Sigma_{\rm e,*},b_{\rm s},R_{\rm e,*},n)\right]^{2}}{M_{i}}\,$, (12)
where $M_{i}$ and $\hat{M}_{i}$ are the stellar masses obtained from the simulation and expected from the fitting function, respectively, in the $i$-th bin of $r_{\rm p}$.
2.1.5 The cold-gas fraction and internal instability
The transitional features in the morphology, dynamical hotness and star-formation burstiness suggest a star-formation environment in the early universe that is substantially different from that in the local universe. To gain insight into the physical origin of the transition, here we follow the ‘quadrant’ diagram introduced by MCW24 (see their §3.3) to consider two factors, one internal and the other external. The first factor, introduced in the following, is related to the cold gas fraction in a galaxy that determines the instability of the gas, while the second, related to the assembly of dark matter halos, is introduced in §2.2. We will use both to understand the evolution of dynamical hotness and transition of galaxy properties in §4.
To quantify the internal instability caused by the self-gravity of a gaseous cloud or a disk against the support of angular momentum, we use the ratio between the cold-gas fraction and the gas spin parameter:
$f_{\lambda}=\frac{f_{\rm gas}}{\lambda_{\rm gas}}\,$, (13)
where $f_{\rm gas}$ is the ratio of the mass of cold gas (with temperature $T<2\times 10^{4}\,{\rm K}$) within $R_{\rm 90,*}$ to the virial mass of the dark matter halo ($M_{\rm vir}$), $\lambda_{\rm gas}=|{\bm{j}}_{\rm gas}|/\left(\sqrt{2}V_{\rm vir}R_{\rm vir}\right)$, with ${\bm{j}}_{\rm gas}$ being the specific angular momentum of the cold gas within $R_{\rm 90,*}$. Fig. 2(f) shows $f_{\rm gas}$ and $\lambda_{\rm gas}$ versus $t$ for m12f. This example shows that the galaxy starts from an early phase where both $f_{\rm gas}$ and $\lambda_{\rm gas}$ are fluctuating around relatively high values. From $z\approx 3$ to $\approx 1.5$, $f_{\rm gas}$ is much higher than $\lambda_{\rm gas}$, and at $z<1$ the order reverses.
The ratio $f_{\lambda}$ is defined by Mo et al. (1998, see their §3) to describe the rotational support of a gaseous disk against self-gravity, with the assumption that the gaseous disk is embedded in a dark matter halo characterized by a static NFW profile that dominates the gravity of all collisionless components. Here we utilize the spin of cold gas, instead of that of dark matter, to account for the potential difference between the two (Jiang et al., 2019; Yang et al., 2023). Depending on the concentration of the halo, the baryon fraction and distribution, the threshold for the gaseous disk to be stable against self-gravity varies around $1$ (e.g. Efstathiou et al., 1982; Mo et al., 1998). Thus, $f_{\lambda}\lesssim 1$ ($\gtrsim 1$) implies a stable (unstable) disk. In §4, we will use the evolution of $f_{\rm gas}$, $\lambda_{\rm gas}$ and $f_{\lambda}$ to understand the formation of different components of galaxies.
The ratio $f_{\lambda}$ defined above is a condition for the gas to become self-gravitating before it can be supported by angular momentum. To study how gas collapses produce a self-gravitating structure near the halo center, we define a parameter,
$f_{\text{gas+star}}=\frac{M_{\rm gas}+M_{\text{young-star}}}{M_{\rm tot}}\,$, (14)
where $M_{\rm gas}$ is the mass of cold gas, $M_{\text{young-star}}$ is the mass of stars formed in the past duration of one dynamical timescale of the host halo, $t_{\mathrm{dyn,h}}(t)\equiv\sqrt{(3\pi)/(16G{\overline{\rho}}_{\rm vir}(t))}$, and $M_{\rm tot}$ is the total mass (dark matter plus baryon mass). All the masses here should be evaluated within a common aperture, which we will specify when we use this parameter. The inclusion of young stars in $f_{\text{gas+star}}$ is to take into account the fact that processes on galactic scales have timescales typically of the dynamical timescale and that some star-forming gas may have already been turned into stars in this duration.
2.2 Dark matter halos hosting the galaxies
The assembly history of a dark matter halo was found to have two phases according to the growth rate (Zhao et al., 2003). Following MCW24 (see their §2), we define the specific growth rate of the virial velocity for a halo at redshift $z$, as
[…]
where $H(z)$ is the Hubble parameter at $z$, $a=1/(1+z)$ is the scale factor, and the differentiation is numerically evaluated along the main branch of the subhalo merger tree. The black solid curve in Fig. 2(g) shows the evolution of $V_{\rm vir}$ for m12f, and that in Fig. 2(h) shows the evolution of $\gamma$ obtained numerically from $V_{\rm vir}$.
The accretion of a halo, as described by $\gamma$, is often noisy, exhibiting significant small-scale fluctuations. Part of the fluctuations may be produced by numerical defects in the halo finder and tree builder. To extract a clean accretion history that perseveres the long-term trend, we follow MCW24 to fit the mass assembly history of each halo by a smooth function of redshift:
$\ln M_{\rm vir}(z)=c_{0}+c_{1}\frac{z}{1+z}+c_{2}\ln(1+z)+c_{3}z\,$, (16)
where ($c_{0},c_{1},c_{2},c_{3}$) are four fitting parameters. The smoothed values of other halo properties, such as virial radius $R_{\rm vir}$, virial velocity $V_{\rm vir}$, and specific growth rate $\gamma$, are derived from the smoothed values of $M_{\rm vir}$. The evolutions of smoothed $V_{\rm vir}$ and $\gamma$ obtained for m12f are shown by the black dashed curves in Fig. 2(g) and (h), respectively. A clear deceleration in the halo accretion, as represented by the evolution of $\gamma$, can be seen. With this smooth fitting, the assembly of a halo can be divided into two phases: an early phase with fast accretion (named ‘fast phase’), and a late phase with slow accretion (named ‘slow phase’). The threshold for such a division falls between a range of $\gamma_{\rm f}\approx 0$$3/8$ (Zhao et al., 2003; More et al., 2015; Boco et al., 2023). Here we use both values to bracket the uncertainty in the division. Specifically, we define transition times, $t_{\gamma}^{\rm(start)}$ and $t_{\gamma}^{\rm(end)}$, as the cosmic times when the smoothed $\gamma(z)$ first reaches the threshold $\gamma_{\rm f}=3/8$ and 0, respectively, and we denote the average of the two transition times as $t_{\gamma}$. In Table 1, we list the transition times for all halos in our sample. The vertical lines in Fig. 2(g) and (h) bracket the range from $t_{\gamma}^{\rm(start)}$ to $t_{\gamma}^{\rm(end)}$ for m12f.
For reference, we also compute the ‘half-mass’ formation time, $t_{\rm h,1/2}$, for each halo in our sample. This formation time is obtained by following the evolution of $M_{\rm vir}$ along the main branch of the subhalo merger tree rooted in the halo at $z=0$, and solving the equation
$M_{\rm vir}(t_{\rm h,1/2})=\frac{1}{2}M_{\rm vir}(z=0)\,.$ (17)
The exact formation redshift is evaluated by linearly interpolating the redshift between the two adjacent snapshots where $M_{\rm vir}$ first crosses the half of its final mass. This time has been commonly used in literature to characterize the assembly of dark matter halos (Gao et al., 2005; Li et al., 2008), and was found to correlate with $t_{\gamma}$ (see Appendix B of MCW24). Another related formation time, which is also commonly used in the literature to characterize halo assembly, is $t_{\rm h,Vpeak}$, defined as the cosmic time when the maximum circular velocity ($V_{\rm max}$) of the halo reaches its peak along the main-branch history. This time was found to be sensitive to violent changes in the halo mass, and has been used to incorporate effects of recent major mergers (Behroozi et al., 2014).
In Table 2, we list the properties of galaxies/halos defined in this subsection and the last. Using these definitions, we obtain the characteristic times for each of the galaxies/halos in our sample, and we list them in Table 1. In general, the piecewise functions introduced in §2.1 fit well the evolution of the properties for most galaxies in our sample. The fitting is less satisfactory for m12w, m12r and m12z. Galaxies m12w and m12z have never reached to the disk-dominated stage, so that the intervals $_{\rm thin}^{\rm(start)}, tt$$_{\rm thin}^{\rm(end)}$] and $t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(start)}$, $t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(end)}$] are wide, and the corresponding transition times, $t_{\rm thin}$ and $t_{\rm\scriptscriptstyle\mathcal{H}}$, are not well constrained. Galaxy m12r has a massive satellite at a distance of $390\,{\rm{kpc}}$ at $z=0$, with stellar mass of $M_{*}\approx 2.8\times 10^{9}{\rm M_{\odot}}$ and a recent pericentric passage at $30\,{\rm{kpc}}$ only about $0.7\,{\rm Gyr}$ ago ($z=0.05$; Samuel et al. 2020). Consequently, the disk of m12r is strongly perturbed, which causes significant fluctuations in $f_{\rm thin}$ and $\mathcal{H}$. The disk of m12r is still preserved during the interaction, primarily because of the largely prograde motion of the satellite with respect to the disk. The exceptional dynamics of these three galaxies also have impact on their burstiness. Galaxies m12w and m12r show significant fluctuations in $\mathcal{B}$ around $0.2$ (the threshold to define $t_{\rm\scriptscriptstyle\mathcal{B}}$), while galaxy m12z never reaches this threshold and so its $t_{\rm\scriptscriptstyle\mathcal{B}}$ is absent in Table 1. Since transitions are quite vague for the three galaxies, we analyze and discuss them separately in the following.
3 Co-evolution pattern between galaxies and halos
[…]
Figure 3: Correlation between halo transition time ($t_{\gamma}$) and galaxy transition times. Each panel in the upper row uses one definition of galaxy transition time: (A), $t_{\rm thin}$; (B), $t_{\rm\scriptscriptstyle\mathcal{H}}$; (C), $t_{\rm\scriptscriptstyle\mathcal{B}}$; (D), $t_{\rm bulge,1/2}$. Each marker represents one galaxy in our sample, with error bar spanning from the start to the end of the transition interval (for $_{\gamma}, tt$$_{\rm thin}$ and $t_{\rm\scriptscriptstyle\mathcal{H}}$) or from 20% to 80% formation time of bulge mass (for $t_{\rm bulge,1/2}$). Markers with grey edges represent the peculiar galaxies (m12w, m12r and m12z) excluded from correlation analysis. For each pair of variables, the Spearman correlation coefficient ($\rho$) is indicated in the corresponding panel (see also Table 3 for a full list of $\rho$ and $p$-values), and black line shows a linear fitting. Each lower panel shows $\Delta T_{\rm dyn}$, the duration of the transition interval corresponding to the upper panel, normalized by the dynamical time of the host halo. This figure demonstrates the strong correlation between halo assembly and galaxy assembly, and implies that the transitions of galaxies are driven by the transition of halo assembly. See §3 for details.
Table 3: Spearman correlation coefficient ($\rho$) and $p$-value between pairs of characteristic times of galaxies and their host halos.
$\rho$ ($p$-value) $t_{\gamma}$ $t_{\rm h,1/2}$ $_{\mathrm{h,Vpeak}} tt_{\mathcal{B}}$
$t_{\mathrm{thin}}$ 0.56 (0.07) 0.30 (0.37) 0.17 (0.61) 0.85 (1e-3)
$t{{}_{\mathcal{H}}}$ 0.70 (0.02) 0.27 (0.42) 0.08 (0.81) 0.85 (1e-3)
$t_{\mathcal{B}}$ 0.82 (2e-3) 0.24 (0.48) 0.30 (0.37) -
$t_{\rm bulge,1/2}$ 0.92 (7e-5) 0.13 (0.69) 0.53 (0.10) 0.8 (3e-3)
In Fig.3(A)–(D), we show how the galaxy transition times ($_{\rm thin}, tt_{\rm\scriptscriptstyle\mathcal{H}}$, $t_{\rm\scriptscriptstyle\mathcal{B}}$ and $t_{\rm bulge,1/2}$) are correlated with the halo transition time ($t_{\gamma}$) using the 14 galaxies in our sample. The solid line in each panel shows the linear fit to the points of individual galaxies, excluding the exceptional cases (m12w, m12r and m12z) identified above. For every definition of the galaxy transition time, a positive correlation with $t_{\gamma}$ is observed: halos with earlier transition to slow accretion on average host galaxies with earlier transition to disk-dominated and dynamically cold phases with lower level of burstiness in star formation. To quantify the significance of these correlations, we compute the Spearman correlation coefficient and the $p$-value between each of the galaxy transition times and the halo transition time, and we list the results in Table 3. The correlation coefficient between $t_{\rm thin}$ and $t_{\gamma}$ is about 0.56, with a $p$-value of 0.07, indicating a significant correlation. The correlations of $_{\rm\scriptscriptstyle\mathcal{H}}, tt_{\rm\scriptscriptstyle\mathcal{B}}$ and $t_{\rm bulge,1/2}$ with $t_{\gamma}$ are even more significant and stronger, as reflected by the higher correlation coefficients of 0.70, 0.82 and 0.92, and the lower $p$-values of $0.02$, $2\times 10^{-3}$, and $7\times 10^{-5}$, respectively. These results suggest that the transitions in galaxy morphology, dynamical hotness and star formation burstiness are closely related to the transition in halo assembly. A similar conclusion was reached by Dillamore et al. (2024) based on the correlation between the ‘spin-up’ time, defined as the time when a galaxy starts to form ordered disk, and the halo assembly for MW-size galaxies in the ARTEMIS simulations.
Besides the strong correlations between galaxy properties and halo assembly, there are significant variations among individual galaxies, indicating that processes other than the halo accretion may also play important roles in determining the transitions in galaxy properties. For example, the galaxy m12m has a $t_{\rm thin}$ about $2\,{\rm Gyr}$ earlier than that expected from the linear fit. A careful examination reveals that m12m has a massive satellite whose orbital motion aligns with the disk rotation of m12m at $z\approx 3$. The transfer of orbital angular momentum, or the transfer of high-circularity stars from the satellite to the disk may thus contribute to the early formation of the thin disk (Jiang et al., 2025). Similar variations were identified by Chen et al. (2021a) and Chen et al. (2021b) by applying a set of successively refined regressors to the galaxies in the IllustrisTNG and EAGLE simulations, where nuanced factors other than halo assembly are inferred to affect star formation in galaxies.
For comparison, the correlations of the galaxy transition times with the half-mass formation time ($t_{\rm h,1/2}$) or the peak-velocity time ($t_{\rm h,Vpeak}$) of the host halo are much weaker, with $p$-values not low enough to reject the null hypothesis of no correlation. A similar result was reported by Garrison-Kimmel et al. (2018), who found a poor correlation between the stellar morphology of galaxies and the half-mass formation time of halos. These findings suggest that it is the transition in the halo accretion rate, rather than the accumulative amount of the accreted mass, that determines the transitions in galaxy properties.
Fig. 3(a) and (b) show the durations of the transitional phases versus the halo transition time, $t_{\gamma}$, for individual galaxies in our sample. The duration of transition for each galaxy is normalized by the dynamical timescale of the host halo, $t_{\rm dyn,h}$, as
[…]
Figure 4: The quadrant diagram showing the evolution of galaxies in the $f_{\lambda}$-$\gamma$ plane. In each panel, dots show the snapshots of all galaxies in our sample, color-coded by a property labeled in the color bar. $f_{\text{gas+star}}$ in (b) are evaluated within the inner $1\,{\rm{kpc}}$ of galaxies. The growth rate of stellar bulge ($\dot{M}_{\rm bulge}$, in $\rm M_{\odot}yr^{-1}$) shown in (h) is smoothed by the LOESS method (Cleveland & Devlin, 1988), using the Python package (Cappellari et al., 2013). White path linking the dots with black edges shows the evolution of m12f. Two vertical lines show $\gamma=3/8$ and $\gamma=0$, respectively, which define the transitional regime of halo assembly. One horizontal line shows $f_{\lambda}=1$, above which self-gravity of gas is expected to prevent the formation of a dynamically cold disk. The $f_{\lambda}$-$\gamma$ plane is partitioned into four quadrants, Q1–Q4, by these lines, within which galaxies are expected to have distinct dynamical states. The FIRE-2 galaxies follow a general evolutionary trajectory in this plane, from Q1 at high $z$ to Q3 at low $z$. See §4.1 for details.
4 Physical processes driving the two-phase growth of galaxies
The correlation pattern between galaxies and halos found above motivates us to identify processes that drive the transition of galaxies based on the halo assembly. According to MCW24 (see their §3.3), two factors, one external and one internal, are predicted to be important in determining the dynamics of galaxies.
The first factor is the assembly rate of the host halo, as quantified by $\gamma$. A halo in its fast phase has a nearly constant concentration of $c_{\rm v}\approx 4$ (Zhao et al., 2003, 2009), reflecting a self-similar growth without changing the shape of the profile. Such an evolution of halo structure relies on a sufficient exchange of energy between the newly accreted matter and the existing matter in the halo. This is feasible only if the accretion is fast enough so that the perturbation of the gravitational potential is frequent and strong. The fast accretion of dark matter also has important implications for baryonic processes: fast accumulation of gas is expected to be associated with fast accretion of dark matter. This, together with the effective radiative cooling expected before halos reach a threshold of $M_{\rm v}\approx 10^{12}{\rm M_{\odot}}$ (see the cooling diagram in, e.g., fig. 2 of MCW24), can cause large amounts of cold gas to fall almost freely into the halo center. Rapid inflow of gas not only sustains a high gas fraction that leads to the formation of a self-gravitating gas cloud (SGC) before angular-momentum support can be established, but also strongly perturbs the SGC and generates supersonic turbulence on the galactic scale. Density fluctuations are seeded by the gravitational perturbations and amplified by shocks in the turbulent SGC, causing the formation of sub-clouds that are dense enough to be free of cloud-cloud collision and drag force from surrounding gas. The almost ballistic motion of sub-clouds makes the mixing of angular momentum inefficient, prevents the formation of a dynamically cold disk and preserves the hot gas dynamics. Star clusters can form within the sub-clouds and inherit the hot dynamics, leading to the formation of a bulge-like stellar system. A dynamically cold gaseous and stellar disk is expected to form only after the halo assembly slows down ($\gamma\lesssim\gamma_{\rm f}$) and the gas fraction is reduced by, e.g. feedback, to a level that is comparable to the gas spin ($f_{\lambda}\equiv f_{\rm gas}/\lambda_{\rm gas}\lesssim 1$). This scenario predicts that the galaxy has hot (velocity-dispersion supported) dynamics at high $z$ and cold (rotation supported) dynamics at low $z$.
The second factor is the relative importance between the self-gravity and angular momentum support of the gas, as quantified by $f_{\lambda}$, the ratio between the cold-gas fraction and the spin. This factor quantifies the internal instability of the disk. A high $f_{\lambda}$ can cause a global instability to develop in the disk, which is expected to thicken or destroy the cold disk even if it can be formed in the first place. Making $f_{\lambda}$ sufficiently small (again $f_{\lambda}\lesssim 1$) is thus a necessary condition for the cold disk to form and to remain for a long period.
4.1 Evolutionary trajectory in the quadrant diagram
The combination of the two factors partitions the evolution of galaxies into four quadrants, Q1–Q4, according to their locations in the $f_{\lambda}$-$\gamma$ plane. The dynamical states of galaxies in different quadrants are predicted to be distinct, as summarized below.
  1. 1.
    Q1: $\gamma\gtrsim\gamma_{\rm f}$ and $f_{\lambda}\gtrsim 1$. This is where all galaxies were born at high $z$. A bulge-like morphology is expected in this quadrant once stars form in the sub-clouds that move randomly. However, the concentration of the bulge is not necessarily high, since stars may form in sub-clouds at large radii and dynamical heating associated with the fast halo assembly and star formation can prevent the build-up of a concentrated structure.
  2. 2.
    Q2: $\gamma\lesssim\gamma_{\rm f}$ and $f_{\lambda}\gtrsim 1$. This is where unstable disks are expected to form. External perturbations from the fast halo assembly are now alleviated, and mixing of angular momentum is elevated, both promoting the formation of a self-gravitating, disk-like structure. However, local instability in the heavy disk can lead to the formation of clumps that can heat the disk dynamically (e.g. Meng & Gnedin, 2020; Romeo et al., 2020; Renaud et al., 2021; van Donkelaar et al., 2022; Romeo et al., 2023). Global instability of the disk can drive rapid gas inflow towards the galactic center to form a bar or pseudo-bulge with a certain level of dynamic hotness (e.g. Du et al., 2020; Zana et al., 2022).
  3. 3.
    Q3: $\gamma\lesssim\gamma_{\rm f}$ and $f_{\lambda}\lesssim 1$. This is the regime in which a cold disk can form, since both external and internal barriers to the coldness are removed. To form a cold stellar disk, $f_{\rm gas}$ cannot be too low, because otherwise the gas cannot cool efficiently to form stars.
  4. 4.
    Q4: $\gamma\gtrsim\gamma_{\rm f}$ and $f_{\lambda}\lesssim 1$. This is the regime where the gas fraction has been reduced by e.g. feedback processes, but dynamical heating by the fast halo assembly is still present. Depending on the structure formed before this stage, the galaxy can be a bulge or thick disk.
Note that the first factor, $\gamma$, is determined solely by the halo assembly and is thus robust regardless of the detailed modeling of the baryonic physics in simulations. However, the second factor, $f_{\lambda}$, depends strongly on baryonic physics, particularly on star formation, black-hole growth, and the feedback processes generated by them. Hydrodynamic simulations with distinct subgrid models can thus predict diverse evolution pathways of galaxies in the $f_{\lambda}$-$\gamma$ plane. This also implies a rich set of possibilities for morphological transformation when galaxies in the simulation migrate from one quadrant to another, thereby providing observational consequences to test models.
[…]
Figs. 4(c)–(g) show the evolution of the burstiness ($\mathcal{B}$), hotness ($\mathcal{H}$), and the fractions of the three kinematic components (bulge, thick disk, and thin disk), respectively. All these quantities exhibit obvious evolution when galaxies migrate from Q1/Q4 to Q3, revealing a general trend that galaxies evolve from a dynamically hot phase, with bulge-like morphology and highly bursty star formation, to a dynamically cold phase, with cold-disk morphology and steady star formation. All these quantities show significant correlations with $\gamma$, and make transitions at $\gamma\approx\gamma_{\rm f}=0$$3/8$. This confirms the findings in §3 that the two phases of galaxy evolution are correlated with the two phases of halo assembly. A weak trend is evident that, given $\gamma$, galaxies with higher $f_{\lambda}$ tend to have a higher hotness and bulge fraction, verifying the expected importance of $f_{\lambda}$ in determining the dynamics of galaxies. All these results support the use of the quadrant diagram to separate the phases of galaxy formation and to identify the physical processes that drive the transitions, as discussed in the following.
Figure 5: Evolution of properties related to the formation of stellar bulges. (a), bulge mass ($M_{\rm bulge}$, normalized by the value at $z=0$) as a function of time. (b), mass fraction of cold-gas plus young stars ($f_{\text{gas+star}}$) within some apertures. Colored markers represent $f_{\text{gas+star}}$ evaluated within the inner $1\,{\rm{kpc}}$ in the history of one galaxy, as labeled in Fig. 3. Black curve bound by grey shade shows the mean and $25^{\rm th}$$75^{\rm th}$ percentiles among all galaxies, with $f_{\text{gas+star}}$ evaluated within the $R_{\rm 50,*}$ of young bulge stars. The time variable of each galaxy is measured with respect to the start time of transition of halo assembly, $t^{\text{(start)}}_{\gamma}$, and normalized by the duration of the transition, $\Delta t_{\gamma}\equiv t^{\text{(end)}}_{\gamma}-t^{\text{(start)}}_{\gamma}$. We highlight the peak value of $f_{\text{gas+star}}$ along the assembly history of each galaxy by a larger marker in (b). For m12r, the peak of $f_{\text{gas+star}}$ is reached at $t-t^{\text{(start)}}_{\gamma}\approx 6\Delta t_{\gamma}$, too late to be shown, so we mark it with an orange arrow. Grey vertical band in each panel shows the transitional regime of halo assembly. See §4.2 for details.
Figure 6: Evolution of galaxy sizes. (a), sizes of the galaxy m12f by different definitions as functions of redshift, pink solid (dashed) for $R_{\rm 50,bulge}$ ($R_{\rm 50,young\,bulge}$), the half-mass radius of all (young) bulge stars, and dark blue for the half-mass radius of all stars formed at $z\geqslant 8$. For comparison, we show the evolution of virial radius, $R_{\rm vir}$ (black), and scale radius $r_{\rm s}\equiv R_{\rm vir}/c$ (grey), of the host halo, where $c$ is the concentration parameter obtained by fitting the radial density profile with the NFW form (Navarro et al., 1996, 1997; Bullock et al., 2001). Pink vertical line indicates $t_{\rm bulge,1/2}$, the half-mass formation time of the stellar bulge. (b), (c), $R_{\rm 50,bulge}$ of all galaxies in our sample, normalized by either the value at $z=0$ (b) or the scale radius of their host halo at the corresponding time (c). The time variable of each galaxy is measured with respect to $t^{\text{(start)}}_{\gamma}$ and normalized by $\Delta t_{\gamma}\equiv t^{\text{(end)}}_{\gamma}-t^{\text{(start)}}_{\gamma}$. Black circles show the mean values among all galaxies. The markers for each galaxy are styled the same as Fig. 3. In each panel, grey vertical band shows the transitional regime of halo assembly. See §4.2 for details.
4.2 The origin of dynamical hotness: concentrated versus scattered star formation
The above analyses based on the quadrant diagram pinpoint two regimes for the build-up of the dynamically hot stellar system in a galaxy. The two regimes are well separated in the $f_{\lambda}$-$\gamma$ plane, and exhibit distinct patterns in the physical properties of galaxies. The later regime is at $\gamma\approx 0$$3/8$, where the fraction of cold-disk stars is about to rise (see e.g. Fig. 2a), but the SFR of the bulge component is the highest among all redshifts in the history of a galaxy (see Fig. 4h). In Fig. 5(a), we show the assembly histories of the bulge mass for all 14 galaxies in our sample. All of these assembly histories have a sigmoid-like shape, with clear transition during the transition phase of halo assembly, $t_{\gamma}^{\rm(start)}\lesssim t\lesssim t_{\gamma}^{\rm(end)}$. Bulge masses grow rapidly in this duration, and most bulge stars at $z=0$ form during this period. The strong correlation between bulge formation time ($t_{\rm bulge,1/2}$) and halo transition time ($t_{\gamma}$) shown in Fig. 3(D) is thus explained. The earlier regime is at $\gamma\gtrsim 3/8$, before the halo assembly slows down. Star formation in this regime contributes only a minor fraction of the final bulge mass at $z=0$ (see Fig. 5a), but almost all stars are formed in dynamically hot components (bulge or thick disk; see Figs. 2c and 4e, f). The two regimes thus have important implications both for observations targeting the old bulge stars at low $z$ and for deep observations targeting the ongoing bulge formation at high $z$, as we will discuss in §5.
The evolutionary pattern of galaxies in the quadrant diagram also provides clues to identify the origin of the hotness in each of the two regimes. One critical feature for the properties of a galaxy in the later regime of bulge formation is that the star-forming gas in the inner region of the galaxy is globally self-gravitating. This is evident from Fig. 4(b), where the fractions of cold gas plus young stars ($f_{\text{gas+star}}$) in the inner $1\,{\rm{kpc}}$ of the galaxies are seen to reach $0.5$ when $\gamma\approx 0$$3/8$, indicating a self-gravitation of the star-forming gas. To see this more clearly for individual galaxies, in Fig. 5(b), we show the evolution of $f_{\text{gas+star}}$ for all the 14 galaxies in our sample, with the time axis aligned to $t_{\gamma}^{\rm(start)}$ and scaled by the duration of the transitional phase, $\Delta t_{\gamma}\equiv t_{\gamma}^{\rm(end)}-t_{\gamma}^{\rm(start)}$. The radius to measure $f_{\text{gas+star}}$ is chosen to be either $R_{\rm 50,young\,bulge}$, the half-stellar-mass radius of young bulge stars (black curve with gray shade) or $1\,{\rm{kpc}}$ (scatter points). Before $t_{\gamma}^{\rm(start)}$, $f_{\text{gas+star}}$ is low in most cases, indicating that no global self-gravitation of the star-forming gas is present. Between $t_{\gamma}^{\rm(start)}$ and $t_{\gamma}^{\rm(end)}$, $f_{\text{gas+star}}$ increases significantly and reaches about $0.5$, indicating that the star-forming gas becomes globally self-gravitating. At $t\gtrsim t_{\gamma}^{\rm(end)}$, $f_{\text{gas+star}}$ within $1\,{\rm{kpc}}$ drops again to a low value, while $f_{\text{gas+star}}$ within $R_{\rm 50,young\,bulge}$ remains close to $0.5$. This suggests that the star-forming gas responsible for the bulge formation is still globally self-gravitating, but its size has shrunk below $1\,{\rm{kpc}}$, a value much smaller than the size of the thin disk component that dominates the star formation at low $z$ (see also Fig. 6 to be discussed below). Bulge formation in this later stage, $t\gtrsim t_{\gamma}^{\rm(end)}$, is highly inefficient and contributes only a small fraction to the instantaneous SFR and the final bulge mass.
The above analyses suggest the origin of hotness in the later regime of bulge formation. Once the spatial distribution of the star-forming gas in a galaxy becomes concentrated enough so that the self-gravitation condition is satisfied, violent fragmentation of the gas naturally arises owing to various types of instabilities, such as the Jeans instability. Sub-clouds formed in this way then produce star clusters, which inherit the hot dynamics of the gas originated from the rapid inflow and the gravitational perturbations associated with the build-up of the high gas concentration. The total mass formed in this way can be very large as a result of the rich amount of gas accumulated during the fast phase of halo assembly and the deepened gravitational potential well of the host halo that resists gas loss by feedback. The formation of the stellar bulge proceeds so rapidly that a dynamically cold gaseous disk does not have enough time to form. Given the importance of high gas concentration in driving the formation of stellar bulges in this regime, we summarize the corresponding origin of the hotness as ‘concentrated star formation’. In Fig. 6(a), we show the evolution of the half-stellar-mass radii of different stellar components of m12f, in comparison to the scale radius ($r_{\rm s}$) and virial radius ($R_{\rm vir}$) of the host halo. Between $t_{\gamma}^{\rm(start)}$ and $t_{\gamma}^{\rm(end)}$, the growth of $R_{\rm vir}$ continues, but the size of the distribution of young bulge stars decreases significantly. The small size of newly formed bulge stars in comparison with the large halo size confirms the scenario of concentrated star formation inferred above from self-gravitation. At $t\gtrsim t_{\gamma}^{\rm(end)}$, the size of newly formed bulge stars drops to a few hundred parsecs, indicating that the concentrated star formation continues to take place, albeit at a very low level and within a very small region. The behavior in the evolution of the bulge size seen in m12f, is also shared by other galaxies, as shown by Fig. 6(b). In all cases, the bulge radius first increases rapidly, reaching a maximum value at $t\approx t_{\gamma}^{({\rm start})}$, and then decreases slowly with time, eventually reaching a roughly constant value at $t>t_{\gamma}^{({\rm end})}$. As shown in Fig. 6(c), when normalized by the scale radius of the host halo at the redshift when $R_{\rm 50,bulge}$ is measured, the ratio is roughly a constant of $\sim 0.3$ at $t<t_{\gamma}^{\rm(start)}$, with large fluctuations generated by the rapid change of the gravitational potential. The ratio then decreases by a factor of about three and settles around $\sim 0.1$ at $t>t_{\gamma}^{\rm(end)}$. As described in Chen et al. (2024), in the two-phase model of galaxy formation, the bulge size is related to $r_{\rm s}$ as $R_{\rm 50,bulge}\approx\alpha r_{\rm s}$, with $\alpha$ being a redshift- and mass-independent constant that is determined by the self-gravitation condition and bulge profile. Matching the prediction of the model with the observational data, Chen et al. (2024) obtained $\alpha\approx 0.04$. As one can see from Fig. 6(c), most of the simulated galaxies follow the model scaling relation after the concentrated formation of the bulge, but their values of $\alpha$ are larger by a factor of more than two. This discrepancy may partly arise from the fact that simulated bulges include all stars in the outer region where the surface brightness may be too low to be account for in the observational data. Another possibility is that the bulge to halo mass radios in simulated galaxies are larger than in real galaxies. Indeed, the stellar to halo mass ratios of FIRE-2 or FIRE-set galaxies (see Table 1; see also Hopkins et al. 2018; Schaye et al. 2025) are more than a factor of two higher than in observations (e.g. Yang et al., 2012).
In the earlier regime of bulge growth, the star-forming gas is not globally self-gravitating, as seen from the low $f_{\text{gas+star}}$ in Fig. 5(b) during $t\lesssim t_{\gamma}^{\rm(start)}$. This, in combination with the high gas fractions in, and the small sizes of halos, suggests that star-forming gas must be distributed in an extended way and that the bulge size-halo size ratio is much higher than that in the later regime of bulge formation. The evolution of sizes in Fig. 6(a) clearly shows this point, where $R_{\rm 50,young\,bulge}/R_{\rm vir}$ remains as high as $0.1$ – a value much higher than 0.010.010.0150.015 observed for bulges in the local Universe (e.g. Kravtsov, 2013; Huang et al., 2017; Zanisi et al., 2020; Chen et al., 2024). These results suggest that the origin of hotness in the earlier regime is that the distribution of star-forming gas is sufficiently extended so that the random motion of gas has not been significantly dissipated until stars are born out of the gas. This is totally different from the situation where the angular momentum between gas elements is well mixed to produce a dynamically cold gaseous disk with ordered rotation before most of that gas is converted into stars. This is, however, similar to the situation in the formation of the brightest cluster galaxies (BCGs) in massive halos at low $z$, where significant fractions of stars were formed ex-situ in satellite galaxies and then accreted to the central galaxies to build their extended stellar outskirts (e.g. Bezanson et al., 2009; Huang et al., 2018; Kluge & Bender, 2023). Given the large extension of the star-forming gas and its importance in determining the hot stellar dynamics, we refer to the origin of the hotness in this regime as ‘scattered star formation’.
As suggested by numerical simulations (e.g. Keres et al., 2005; Dekel & Birnboim, 2006; Dekel et al., 2009; Danovich et al., 2015), galaxies at high $z$ are often fed by cold streams of gas that are seeded by large-scale dark-matter filaments and condensed by radiative cooling. Scattered star formation is thus expected to take place in these cold streams. This is indeed the case for FIRE-2 galaxies, as demonstrated by the example shown in Fig. 7(c). The environment provided by a fast-accreting halo at high $z$ can thus be understood as to provide a condition for the IGM/CGM similar to that satisfied in satellite galaxies in massive halos at low $z$, where gas density is sufficiently high to allow local self-gravitation and star formation. Note that the large extents and filamentary structures of newly formed stars also predict features in the size and morphology of high-$z$ galaxies directly testable by deep observations, as we will discuss in §5.
Given the two distinct modes of star formation that build up the dynamically hot stellar component, a question arises as what drives the transition of galaxies from the scattered star formation to the concentrated star formation. This again can be inferred from the evolution of galaxies in the quadrant diagram. An obvious difference between the two regimes is the assembly rates of the host halos. As suggested by N-body simulations (Borzyszkowski et al., 2017), fast-accreting halos are often connected with filaments in the cosmic web, while slow-accreting halos are not. The cold streams feeding the formation of bulges in the earlier regime ($\gamma\gtrsim 3/8$) are thus expected to disappear in the later regime ($\gamma\approx 0$$3/8$), as dark matter filaments seeding the gaseous streams disappear. Consequently, the high density condition for scattered star formation in CGM/IGM vanishes in the later regime, allowing the gas to accumulate into the halo center without being consumed by star formation and feedback until the condition of self-gravitation is satisfied globally to trigger an intensive and concentrated star formation. During the transitional phase, the potential well of the host halo, as quantified by the maximum circular velocity, $V_{\rm max}$, has already reached the peak value among all redshifts (Zhao et al., 2003). This, together with the rich amount of gas accumulated during the fast phase and the self-gravity of the gas that further deepens the potential well, creates a condition that is highly resistant to feedback. A large amount of gas can thus be converted into stars in a short period, which explains the rapid growth of bulge mass during the concentrated star formation.
Another mechanism that may also contribute to the transition from the scattered to the concentrated star formation is ‘dynamical heating’, by which stars formed at early epochs are ‘heated’ later on by gravitational perturbations. One source of dynamical heating is the fast assembly of the dark matter halo, whose effect is expected to be strong at high $z$. The large extents of bulges formed via scattered star formation at high $z$ imply that the total gravity on the bulge is dominated by that of the dark matter. As halos in the fast phase self-similarly expand its structure (Wechsler et al., 2002; Zhao et al., 2003), stars co-expand with dark matter as they gain energy from the fast-varying gravitational potential. In Fig. 6(a), we show the evolution of the half-mass radius of the stars formed at $z\geqslant 8$ (dark-blue curve). The half-mass radius of these old stars expands with time, suggesting that dynamical heating is indeed effective. The long-term trend of such an expansion follows that of the halo radius, indicating that dynamical heating by fast halo assembly is likely an important heating mechanism. The other sources of dynamical heating can be from episodic outflow produced by bursty star formation that causes temporary loss of binding energy of stars, and the formation of dense gas/stellar clumps that scatter stars and transfer energy to them. Such effects have been suggested by e.g. El-Badry et al. (2016), who showed that bursty star formation leads to outward migrations of stars, more so for older stars (Kado-Fong et al., 2022), and by McCluskey et al. (2024), who demonstrated that the velocity dispersion of stars rapidly increases with time at high $z$. Separating the sources of dynamical heating, however, is not straightforward, as both dark-matter and baryonic sources of dynamical heating are expected to be more effective at higher $z$. Future studies utilizing idealized simulations can help to disentangle these sources, and to quantify their relative importance.
While the transition between the two regimes of bulge formation is clearly seen in our sample of galaxies, we note that this largely relies on the specific sub-grid models adopted in the FIRE-2 simulations: once gas is accreted into a halo through filaments and condensed to a density sufficient for star formation, the ‘early’ feedback produced by young stars through, e.g. stellar winds and ionizing radiation, may be able to disperse the gas before more stars can form. Such feedback from star formation has been investigated in some hydrodynamic simulations for isolated galaxies (Übler et al., 2014; Agertz & Kravtsov, 2015, 2016; Li et al., 2020; Deng et al., 2024), and is expected to be important also in cold filaments due to their shallower gravitational potential well than that in the ISM. In FIRE-2, the early feedback was included in the form of internal radiative feedback from the stars and external radiative feedback from the UV background, but the implemented feedback strength may not be sufficient to prevent star formation in the filaments. Future observations able to detect the low-surface-brightness outskirts of high-$z$ galaxies are crucial to test the presence of the scattered star formation and to narrow down sub-grid models to be implemented in simulations.
4.3 Processes driving the transition to cold disk
As discussed above in connection to the quadrant diagram, the formation of a thin disk requires two conditions. First, the gravitational potential well should be stable so that dynamical heating by variation of the gravitational potential is not important. A high accretion rate of the halo, represented by a high value of $\gamma$, is a primary driver of dynamical heating on galactic scales. Thus, the formation of thin disks is expected only when $\gamma$ is low, as is demonstrated clearly by the strong correlation between $t_{\rm thin}$ and $t_{\gamma}$ in Fig. 3. The second condition is that the cold gas fraction must be sufficiently low so that the total gravitational potential in the inner region of the halo is not dominated by the disk. This is demonstrated by the correlation between $f_{\rm thin}$ and $f_{\lambda}$ shown in Fig. 4(g), and by the rapid decline of $f_{\text{gas+star}}$ within inner $1\,{\rm{kpc}}$ at $t>t_{\gamma}^{({\rm end})}$, shown in Fig. 5(b), when the star formation starts to be dominated by that in the thin disk (see Fig. 2).
In the simulated galaxies, the gravitationally dominant component undergoes two transitions. Initially, at $t<t_{\gamma}^{({\rm start})}$ when $\gamma$ is large, dark matter dominates the gravitational potential almost everywhere in the halo, as implied by the large sizes of galaxies with respect to the sizes of halos (see Fig. 6c). During this phase, a small amount of stars is formed in the ‘scattered mode’ discussed in §4.2, and the gravitational potential near the halo center is still dominated by dark matter. However, as the evolution moves to the transitional phase of halo assembly, $t_{\gamma}^{({\rm start})}<t<t_{\gamma}^{({\rm end})}$, a large amount of low-spin gas accumulates in the halo center, becomes self-gravitating, forms bulge stars in the ‘concentrated mode’, and reduce the gas by feedback. As shown in Fig. 5(b), in almost all simulated galaxies, the gravitational potential in the central part of the halo becomes dominated by star-forming gas and young stars in the bulge in this phase. As soon as the bulge component is established at $t>t_{\rm bulge,1/2}\sim t_{\gamma}$, it provides another source to stabilize the gravitational potential and to promote the formation of a dynamically cold disk from gas with sufficiently high spin (see also Hopkins et al., 2023). As the disk grows, the bulge component can continue to grow slowly from low-spin gas near the center, as shown in Fig. 1f. Thus, the transition to thin-disk formation in the simulated galaxies is well described by the moving of a galaxy to Q3 in the quadrant diagram. In particular, the formation of a dynamically hot bulge in the fast accretion regime of the halo promotes the formation of a thin disk by stabilizing the gravitational potential as well as by reducing the amount of low-spin gas, making the transition to thin-disk formation at $t>t_{\gamma}^{({\rm end})}$ more prominent.
Figure 7: Evolution of the structure of the stellar bulge of m12f. (a)(c), 2-D stellar mass surface density of the bulge component at three redshifts, respectively. Dashed circle in each panel indicates $R_{\rm 50,bulge}$. (d), stellar mass surface density as a function of galactocentric distance ($r$), evaluated using bulge stars from the simulation (solid) and fitted with a Sérsic profile (dashed) at three redshifts. For comparison, starred markers show the profile using disk (thin + thick) stars. (e), Sérsic index of the stellar bulge as a function of cosmic time. (f), minor-to-major axis ratio ($c/a$; solid) and triaxiality ($T$; dashed) of the stellar bulge as functions of cosmic time. In (e) and (f), vertical red solid (dashed) line represents $t_{\gamma}$ ($t_{\mathcal{H}}$). Starred markers indicate the three redshifts displayed in (a)(d), and the corresponding structural parameters are labeled in (a)(c). In (e), gray vertical band marks the range of redshift where the Sérsic fitting is ill-defined due to the temporal and spatial fluctuations in the surface density profiles. This figure highlights that bulge evolves from a prolate, clumpy and low-concentration structure at high $z$ to a spherical, smooth and high-concentration structure at low $z$. See §5 for details.
Figure 8: Evolution of structural parameters of stellar bulges. (a)(d), evolution of the Sérsic index ($n$), triaxiality ($T$), and axis ratios $c/a$ and $b/a$, respectively, of the stellar bulges (red; curve for the mean and shade for the $25^{\rm th}$$75^{\rm th}$ percentiles among all galaxies in our sample). For comparison, we also show the same parameters for all stars (blue). The time variable of each galaxy is measured with respect to $t^{(\rm start)}_{\gamma}$ and normalized by $\Delta t_{\gamma}\equiv t^{(\rm end)}_{\gamma}-t^{(\rm start)}_{\gamma}$. Grey vertical band in each panel indicates the transitional regime of halo assembly. (e), the migration of bulges in the $c/a$-$b/a$ plane, with each pair of empty and filled markers linked by a line representing the bulge of a galaxy in the hot and cold phases, respectively. The evolution of bulge structure implies the two distinct modes of star formation that drive the dynamical hotness, which may be tested by JWST and future observations. See §5 for details.
Figure 9: Distribution of stellar bulges in the $b/a$-$c/a$ plane. Here we only include young bulge stars in given ranges of galactocentric distance: (a), $r<0.25R_{\rm 90,*}$; (b), $0.25R_{\rm 90,*}<r<0.5R_{\rm 90,*}$. Data points show the snapshots of all galaxies in our sample, with color coded according to the redshift smoothed by the LOESS method (Cleveland & Devlin, 1988). In each panel, squares show the evolution of the axis ratios obtained by the median of all galaxies. This figure verifies that scattered star formation in high-$z$ galaxies takes place at large radii in cold streams, thus producing anisotropic shapes of bulges. See §5 for details.
5 Structural evolution of dynamically hot stellar systems
A particularly interesting part of the two-phase formation of galaxies is the growth of dynamically hot stellar systems (referred to as bulges in this paper) in fast-accreting halos. The importance of bulge formation is threefold. First, the ‘scattered formation’ of stars in the early epoch depends on the competition between various cooling and heating processes. Many of these processes are unresolved in current simulations and are implemented as subgrid models. The bulge components formed through this channel may thus encode information about these processes, and their properties may be used to distinguish models. Second, as shown in Fig. 4(e) and other previous studies using different simulations (Pillepich et al., 2019; Yu et al., 2023), the bulge components dominate the total stellar mass of high-$z$ galaxies, and thus their properties have important implications for observations targeting high-$z$ galaxies. Recent observations with resolved stellar dynamics reveal surprising cases of disk-like galaxies at $z\gtrsim 4$, which appears at odd with theoretical expectations. Third, bulges are found to co-evolve with the central supermassive black holes (SMBHs) of galaxies (Kormendy & Ho, 2013; Greene et al., 2020; Zhuang & Ho, 2023), and thus closely related to understanding galaxy quenching (Dekel & Burkert, 2014; Woo et al., 2015; Lin et al., 2019; Hong et al., 2023; Mo et al., 2024). The bulges in FIRE-2 may thus be used to study potential sites of SMBH growth. Here we focus on bulges in FIRE-2 by identifying properties that not only reflect the physical processes driving the formation and evolution, but also can be measured observationally.
Within the two-phase model, Chen et al. (2025b) showed that the fraction of star formation in dense star clusters is expected to increase towards high $z$ (Chen et al., 2025b, see their fig. 18), and such a trend can be used to explain the formation of old globular clusters. To see if this expectation is supported by FIRE-2 galaxies, we show, in Fig. 7(a)–(c), the projected stellar distribution of the bulge component of the galaxy m12f at three redshifts chosen to represent different stages of bulge formation. The change in the clumpiness with time is obvious: the galaxy is highly clumpy at $z=4.4$, becomes less clumpy at $z=1.5$, and appears smooth at $z=0$. A similar finding was also reported by Ma et al. (2020) for a number of other galaxies in the FIRE-2 suite.
Aided by strong lensing, recent observations by JWST revealed an excess of high-$z$ galaxies that are dominated by dense stellar clumps (e.g. Vanzella et al., 2023; Lin et al., 2023; Mowla et al., 2024; Adamo et al., 2024; Fujimoto et al., 2025), supporting the expectation that stellar dynamics is hot in high-$z$ galaxies. Unfortunately, most galaxies without being strongly lensed are not well resolved, and so it is difficult to make a conclusion for the overall galaxy population. Even more challenging is to quantify potential selection bias in such observations. Similar limitations in resolution and selection exist for observations that depend on resolved gas dynamics (e.g. Parlanti et al., 2023; Rowland et al., 2024), as the selection is naturally biased toward galaxies with large masses and significant rotation.
Given the different formation channels of dynamical hotness found in §4.2, we expect that the structural properties of bulges should evolve with time. Bulge sizes, which have been discussed in §4.2 and Fig. 6, are expected to be large during the scattered formation at the fast accretion phase, and to stop growing or even shrink during the later concentrated formation. Bulges at high $z$ are also expected to show significant elongation in shape and small concentration in radial profile. The expectation of elongated shapes for high-$z$ galaxies is based on the scattered star formation in dense gas clumps associated with cold streams (e.g. Mandelker et al., 2018; Danovich et al., 2015). Stars formed in such a way are expected to inherit the overall velocity field of dominating streams, making the stellar distribution preferentially along the major axis of the velocity ellipsoid. The low concentration is expected directly from the ‘scattered’ nature of star formation together with subsequent dynamical heating that makes the stellar distribution more extended. In contrast, bulge formation at lower $z$ is mainly fed by the hot-mode accretion that is more isotropic, with larger amounts of dissipation of binding energy in the halo. Dynamical heating is also less significant at low $z$, all together making the bulges more spheroidal and more concentrated.
The evolution of the bulge structure can be seen from the example images shown in Fig. 7(a)–(c). At $z=4.4$, the stellar distribution is elongated along the diagonal direction of the image, exhibiting filament-like structures that trace the accretion flow of gas into the galaxy. We visually inspected the cold gas stream and found that it is indeed aligned with the stellar distribution. Meanwhile, the stellar distribution is extended. The size of the bulge (indicated by a dashed circle; see also Fig. 6a) is already comparable to the size at $z=0$, although the stellar mass is much lower than the final bulge mass (see Fig. 2a). The concentration of stars in the galaxy is evident, but the multiple dense stellar clumps near the center make it difficult to identify the center. At $z=1.5$, the stellar distribution is less elongated, with a well-defined concentration. At $z=0$, the bulge looks spherical, and the concentrated core is more prominent. In panels (d)–(f), we show the evolution of the stellar mass surface density profile $\Sigma_{*}(r_{\rm p})$], the Sérsic index ($n$), and the shape parameters (axis ratio, $c/a$, and the triaxiality, $T$), respectively, of this bulge. As can be seen, the bulge evolves from a prolate, low-concentration structure at high $z$ to a spheroidal, high-concentration structure at low $z$. There appears to be a clear transition in the evolution of the structural parameters, and the transition time corresponds well to that in halo assembly ($t_{\gamma}$, marked by the red solid lines in panels e and f) and in dynamical hotness of the galaxy ($t_{\cal H}$, marked by the red dashed lines in panels e and f). This provides further support that the evolution of bulges is driven by the evolution of dark matter halos.
In Fig. 8, we show the evolution of the bulge structure for all the 14 galaxies in our example. In panels (a)–(d), we show the evolution of the structural parameters: Sérsic index $n$, triaxiality $T$, $b/a$ and $c/a$, respectively, as functions of cosmic time for the median and $25^{\rm th}$$75^{\rm th}$ percentiles among all bulges. All these results confirm the expectation that the bulge structure evolves from being prolate (small $c/a$ and $b/a$; large $T$) to being spheroidal (large $c/a$ and $b/a$; small $T$), and the bulge concentration (Sérsic index $n$) increases with time. The scatter of $n$ is large at high $z$, reflecting the presence of violent dynamical disturbances that can distort the distribution of stars and prevent the formation of a well-defined galactic center (see also Fig. 7c and f). We checked the Sérsic fitting and found that the $\chi^{2}$ is significantly enlarged at $z\gtrsim 2$, again indicating the distorted stellar distribution and the difficulty of using a single analytical profile to capture such distortion. The cosmic times in panels (a)–(d) are aligned with $t_{\gamma}^{\rm(start)}$ for individual galaxies. Clearly, the transition in shape properties is closely related to the transition of the halo assembly. This tight relation between the transition in bulge shape and the transition in halo assembly seems to indicate that the accretion rate and accretion configuration are correlated with each other, so that the change from fast to slow accretion of the halo corresponds to the change from an anisotropic cold-mode accretion to an isotropic hot-mode accretion of gas into the galaxy. In panel (e), we show the locations of bulges in the c/ac/ab/ab/a plane at two epochs: one is the earliest time when a reliable measurement of the shape parameters can be made for a galaxy (when the number of the star particles is larger than 1000) in the dynamically hot phase, and the other is at $z=0$, when the halo is already in slow accretion. We see clearly that the evolution is from being prolate to being spheroidal (see also Meng & Gnedin, 2021; Xia et al., 2025).
[…]
In observations, resolving structural properties such as the shape and concentration is less demanding than resolving small-scale clumpiness. The evolution of the bulge structure thus provides a promising way to test the two-phase formation model by observations. Indeed, recent observations of high-$z$ galaxies with JWST indicate an excessively large fraction of galaxies with prolate stellar distribution (Law et al., 2012; Yuma et al., 2012; van der Wel et al., 2014b; Zhang et al., 2019; Pandya et al., 2024). However, the small Sérsic index shown in Fig. 8(a) for FIRE-2 galaxies at high $z$ indicates that the measurements of the Sérsic index may not be able to distinguish prolate ellipsoids from inclined disks. Caution is thus needed when interpreting the density profiles of high-$z$ galaxies. For comparison, we show the evolution of the median of individual structural parameters using all stars in Fig. 8(a)–(d). Except for the Sérsic index, the evolution of the structural parameters is less significant than that obtained using only bulge stars, suggesting that a precise decomposition between the bulge and disk components is needed to constrain the formation mechanism of dynamically hot stellar systems.
6 Summary and discussion
In this paper, we use 14 MW-size galaxies from the FIRE-2 cosmological zoom-in simulations to study the co-evolution between galaxies and their host dark matter halos. We focus on the transition behavior of galactic dynamics that separates the evolution into two distinct phases, and highlight the critical role of the two-phase formation of halos in driving the two-phase formation of galaxies. We particularly identify the processes from which dynamical hotness of galaxies originates, and we study the implications of our results for testing these processes using observations. Our main results and conclusions are summarized as follows.
  1. 1.
    We define a set of parameters that either quantify the dynamical state of a galaxy or reflect the consequence of a given dynamical state. These include the stellar masses in different kinematic components (thin disk, thick disk and bulge), the mass fractions of these components in the total stellar mass of a galaxy ($f_{\rm thin}$, $f_{\rm thick}$ and $f_{\rm bulge}$), the dynamical hotness ($\mathcal{H}$), and the burstiness of star formation ($\mathcal{B}$). We find that all these parameters follow a two-phase pattern in their time evolution, and we identify the characteristic times for the evolution of these parameters: the half-mass formation time of the bulge ($t_{\rm bulge,1/2}$), the transition time when mass fraction of the thin-disk component increases rapidly with time ($t_{\rm thin}$), the transition time in dynamical hotness ($t_{\cal H}$) of a galaxy, and the transition time in the burstiness of star formation ($t_{\cal B}$) (§2.1; Figs. 1 and 2).
  2. 2.
    We investigate the co-evolution between galaxies and their halos by linking each of the above characteristic times of galaxies to the time at which the assembly of halo transits from the fast phase to the slow phase ($t_{\gamma}$), and we find that all the characteristic times of galaxies correlate strongly with $t_{\gamma}$. This motivates us to explore the causal connection between the two-phase formation of halos and that of galaxies. The transition of $f_{\rm thin}$ and $\mathcal{H}$ usually completes within 1133 dynamical timescales of the host halo, with a few outliers that take longer time to complete the dynamical adjustment. (§3; Fig. 3).
  3. 3.
[…]
  4. 4.
    The dynamical hotness of galaxies originates from two distinct modes of star formation. The first one is a scattered mode, occurring at $\gamma\gtrsim\gamma_{\rm f}$, in which stars form out of locally unstable gas clumps at large radii in the gaseous streams seeded by the dark matter filaments of the cosmic web. The second one is a concentrated mode, occurring at $\gamma\approx\gamma_{\rm f}$, in which stars form out of sub-clouds fragmented from the entire self-gravitating gas cloud (SGC). The large amount and high density of the accumulated gas, and the deepened gravitational potential well, make the star formation in the concentrated mode highly violent and lead to the formation of the majority ($\sim 70\%$) of bulge stars seen at $z=0$. The migration of star formation from the scattered mode to the concentrated mode is inferred to be caused by the disappearance of gaseous streams and the cease of dynamical heating in slow-accreting halos (§4.2; Figs. 5 and 6).
  5. 5.
    At $\gamma<\gamma_{\rm f}$, the perturbation of gravitational potential by halo accretion ceases. Concentrated star formation has significantly reduced the fraction of gas via violent feedback, and built up a concentrated gravitational potential dominated by the bulge stars, both preventing the gas from self-gravitating globally. Dynamically cold gaseous and stellar disk now have enough time to form. The growth of the bulge via the concentrated mode continues, but is limited to the innermost region where self-gravitation of gas can still be achieved (§4.3).
  6. 6.
    The two modes of star formation may leave observable imprints on the structures of bulges. Bulge stars formed in the scattered mode have large spatial extents (a few tens percent of $R_{\rm vir}$) and small concentration, and are distributed in a prolate shape elongated along the inflow stream. In contrast, bulge stars formed in the concentrated mode have small spatial extents (a few percent of $R_{\rm vir}$) and large concentration, and are distributed in a spheroidal shape. Other observable signatures related to galactic dynamics, such as the burstiness of star formation, and the clumpiness of star-forming gas and young stars, may also be used to decode the mechanisms driving the dynamical evolution of galaxies (§5; Figs. 7, 8 and 9).
The findings here support the two-phase formation scenario of galaxies and reveal the critical role of halo assembly in determining the two phases. However, FIRE-2 is specilized by its subgrid physics model, leading to lack of the compact, baryonic-dominated bulges of the high redshift dwarfs (Shen et al., 2024) and too dispersion- dominated galaxies for the at intermediate masses ($M_{\rm*}\sim 10^{\text{8--10}}\rm M_{\odot}$) (El-Badry et al., 2018; Kado-Fong et al., 2022). Some of our conclusions may depend on the sub-grid models adopted in FIRE-2, For example, as noted in the end of §4.2, the ‘early’ feedback produced by young stars and the external feedback from the UV background may be able to disperse the gas in cold filaments, thereby reducing the amount of stars that can form in the scattered mode (spurious heating term at $z\gtrsim 8$, see e.g. Su et al., 2018; Garrison-Kimmel et al., 2019b). Numerical resolution also matters. The formation of dense gaseous and stellar clumps from the turbulent medium associated with the fast assembly of halos may continuously interact with the surrounding gas and stars, causing significant dynamical heating that limits star formation in concentrated mode (e.g. Mo & Mao, 2004; Ogiya & Nagai, 2022; McCluskey et al., 2024). Failing to properly resolve such clumps may therefore lead to incorrect results. Another issue arises from the specific sample of galaxies selected by the zoom-in runs of FIRE-2, which imposes isolation criteria on the target halos that can bias the sample against dense environments. In Appendix A, we show the distribution of galaxies in the $t_{\rm h,1/2}$-local density plane, and we find that the bias in $t_{\rm h,1/2}$ is small, but that in local density may be significant. Clearly, all these issues need to be studied further using simulations with improved sub-grid models and samples with high completeness.
Another important component that is absent in our analysis is supermassive black holes (SMBHs). As suggested by the analytical estimation in Chen et al. (2025a, see their §4.3 and §4.4), the turbulence of gas plays a fundamental role in breeding the seeds of SMBHs, promoting their early growth, and determining the nuclear structure of galaxies. The growth of SMBHs is thus expected to be closely related to the build-up of the bulge components of galaxies in their hot phase. The rapid growth of bulge during the transitional phase through the concentrated mode coincides with the ‘blue nugget’ phase found in observations (Barro et al., 2014, 2016; Ji & Giavalisco, 2023), which was inferred to be closely related to the boost of AGN activity (Carr et al., 2024; Cohn et al., 2025; Carr et al., 2025; Chen et al., 2025a). The co-evolution among SMBHs, galaxies, and halos should thus be investigated coherently to form a complete picture of galaxy formation. We will return to such an investigation using hydro simulations that include all these components.
Our results have important implications for observations to search for the evidence of scattered star formation using structural properties (size, axis-ratios, triaxiality and Sérsic index) of bulges. Current observations with JWST are already capable of measuring some of these properties for high-$z$ galaxies (Law et al., 2012; Yuma et al., 2012; van der Wel et al., 2014b; Zhang et al., 2019; Pandya et al., 2024). However, because of the flux limit, the measurements may miss the faint outskirts of galaxies that are expected to carry the most significant imprints of anisotropic star formation (Fig. 9). A stacked analysis of a large sample of galaxies may thus be necessary to reveal the full extents of high-$z$ galaxies and to put constraints on the formation mechanisms of galaxies in the early Universe.
Acknowledgements
QM would like to acknowledge Cheng Li and Tao Jing for discussion. YC is funded by the China Postdoctoral Science Foundation (Grant No. 2022TQ0329), and thanks Hui Li and Kai Wang for their valuable insights and suggestions. The authors would like to express their gratitude to the Tsinghua Astrophysics High-Performance Computing platform at Tsinghua University for providing the necessary computational and data storage resources that have significantly contributed to the research results presented in this paper, and to Yi Mao and Chen Chen for their invaluable technical support.
Data Availability
Data and codes used to produce the results in this paper are available upon reasonable request to the corresponding authors. Public data of the FIRE project is available at http://flathub.flatironinstitute.org/fire.
Appendix A The sample of FIRE-2 galaxies
Figure 10: Distribution of FIRE-2 galaxies in the assembly-environment planes. Each colored marker shows the location of a FIRE-2 galaxy at $z=0$ in (a), the $t_{\rm h,1/2}$$\rho_{1}/\bar{\rho}$ plane; (b), the $t_{\gamma}$$\rho_{1}/\bar{\rho}$ plane. Here, $\rho_{1}/\bar{\rho}$ is the local density contrast within $1\,{\rm Mpc}$ around each galaxy. In each panel, black contours, from inner to outer, encompass 25%, 50%, 75% and 90% of halos with masses comparable to the FIRE-2 hosts, taken from TNG300-1-Dark, a dark-matter-only simulation covering a large volume. The FIRE-2 galaxies are not significantly biased in terms of halo assembly at given local density, but those in dense environments are missed. See §A for details.
As briefly introduced in §2, the sample of FIRE-2 galaxies used in this study is drawn from a cosmological simulation, with selection criteria applied to ensure the isolation or the mimicking of the Local Group environment. A question thus arises as to whether our conclusions on the co-evolution between galaxies and halos are biased by the sample selection. Here we address this question by comparing the distribution of our sample with a complete sample drawn from a large-volume cosmological simulation.
For this purpose, we use TNG300-1-Dark, a dark-matter-only simulation performed as a part of the IllustrisTNG project (Pillepich et al., 2018; Nelson et al., 2019). This simulation has a box size of about $300\,{\rm Mpc}$ on a side, and a mass resolution that can resolve halos with $M_{\rm vir}\gtrsim 2\times 10^{9}{\rm M_{\odot}}$. From this simulation, we select a ‘reference sample’ that includes all halos at $z=0$ with $M_{\rm vir}$ in the range of $10^{12}$$10^{12.05}\,h^{-1}{\rm M_{\odot}}$, comparable to the MW-size halos hosting the FIRE-2 galaxies. For halos in FIRE-2 and the reference sample, we calculate $\rho_{1}$, the local matter density within a sphere with a radius of $1\,{\rm Mpc}$ centered on each halo, and normalize it by the mean density of the universe at $z=0$ to get the local density contrast, $\rho_{1}/\bar{\rho}$. We also calculate the half-mass formation time, $t_{\rm h,1/2}$, and transition time, $t_{\gamma}$, for each halo, following the definitions in §2.2. We show the distribution of FIRE-2 halos and TNG300-1-Dark halos in the $t_{\rm h,1/2}$-$\rho_{1}/\bar{\rho}$ and $t_{\gamma}$-$\rho_{1}/\bar{\rho}$ planes in Fig. 10.
All the FIRE-2 halos are found to be embedded in environments with local density contrast $\rho_{1}/\bar{\rho}\lesssim 10$, while the reference sample extends to much higher densities. This suggests that the selection in FIRE-2 is indeed biased towards under-dense environments, as is expected given the isolation criteria used in the simulation. However, at a given local density, the distributions of $t_{\rm h,1/2}$ and $t_{\gamma}$ do not show significant differences between the FIRE-2 sample and the reference sample, implying that the selection does not significantly bias the sample in terms of halo assembly. The analyses in this paper thus fairly represent halos that are sufficiently isolated, while halos susceptible to strong environmental effects due to the proximity to massive cosmic structures are missed.